Solving equations by factoring:
The equation is (2x-3)(3x-8)=0

Answer by stanbon(75887)

Solving equations by factoring:
The equation is (2x-3)(3x-8)=0
-----------
Since the product is zero, one of those factors must be zero.
2x-3 = 0 or 3x-8 = 0
x = 3/2 or x = 8/3
--------------------------
Cheers,
Stan H.
----------------